OpenStudy (anonymous):

lim as x=>3- ((sqrtx^2+16)-5)/(x^2-9)

OpenStudy (anonymous):

rationalise the numerator by multiplying both numerator and denominator by sqrt(x^2+16) +5 you get ( x^2 +16 - 25)/((x^2-9)(sqrt(x^2+16)+5)) hence you get x^2-9 in both numerator and denominator so the limit simplifies to 1/(sqrt(x^2+16)+5) now x=3 so the limit is 1/10.

OpenStudy (anonymous):

sorry, the x^2-9 is also in a sqrt

OpenStudy (anonymous):

ok could you please simplify the question what abt the -5 and the +16?

OpenStudy (anonymous):

ok it's \[(\sqrt{x^2+16}-5)/\sqrt{x^2-9}\]

OpenStudy (anonymous):

Then the limit is just 0